# ETGL-DDPG: A Deep Deterministic Policy Gradient Algorithm for Sparse Reward Continuous Control

### v2

@@ A.1 $\epsilon t$-greedy Sample Efficiency : Proofs
We aim to show that the $\epsilon t$-greedy algorithm falls into the PAC-MDP category. According to Liu & Brunskill (2018), an algorithm $\mathcal{A}$ is PAC-MDP if the covering time induced by $\mathcal{A}$ is polynomially bounded. In Liu & Brunskill (2018), the authors further demonstrate that bounding the covering time requires bounding both the Laplacian eigenvalues and the stationary distribution over the states induced by the random walk policy. This is presented as Proposition A.1. According to Theorem 2, two conditions are satisfied: $N\leq\Theta(|\mathcal{S}||\mathcal{A}|)$ and a lower bound on the probability of the sampled option, $\mathcal{P}_{\mathcal{X}}\geq\frac{1}{\Theta(|\mathcal{S}||\mathcal{A}|)}$. These two conditions are necessary and are met by our problem setting and the exploration algorithm (Algorithm 1). To prove that $\mathcal{P}_{\mathcal{X}}\geq\frac{1}{\Theta(|\mathcal{S}||\mathcal{A}|)}$, we construct a worst-case tree structure $\mathcal{X}$, where we aim to identify the option induced by the tree $\mathcal{X}$ with the lowest probability, referred to informally as the “hardest option". We then show that this lower bound satisfies the condition specified in Theorem 1.
@@
$\mathcal{P}[\text{option returned from }s_{\text{root}}\text{ to }s_{N}\mid\mathcal{E}_{1}\text{ and }\mathcal{E}_{2}]$ $=\prod_{i=2}^{N}\mathcal{P}[\text{(State }s_{i}\text{ added to tree }\mathcal{X})\wedge(s_{i}=s_{min})\wedge(s_{x}=s_{i-1}\text{ in Line 8})]$
@@
[…]
@@
First note that if $\mathcal{P}_{\mathcal{X}}[\omega]\geq\frac{1}{\Theta(|\mathcal{S}||\mathcal{A}|)}$ then based on Theorem 1 we need to have $N\leq\frac{\log(|\mathcal{S}||{\mathcal{A}}|)}{\log\log(|\mathcal{S}||{\mathcal{A}}|)}$, and this implies that $N\leq\Theta(|{\mathcal{S}}||{\mathcal{A}}|)$. Based on the paper by (Liu & Brunskill, 2018), and the analysis of the covering length when following a random policy, we have the following preposition:
Preposition A.1 (Liu & Brunskill (2018)).
: For any irreducable MDP M, we define $P_{\pi_{RW}}$ as a transition matrix induced by random walk policy $\pi_{RW}$ over $M$ and $L(P_{\pi_{RW}})$ is denoted as the Laplacian of this transition matrix. Suppose $\lambda$ is the smallest non-zero eigenvalue of $L$ and $\Psi(s)$ is the stationary distribution over states which is induced by random walk policy, then Q-learning with random walk exploration is a PAC RL algorithm if: $\frac{1}{\lambda}$,$\frac{1}{\min_{s}\Psi(s)}$ are Poly($|{\mathcal{S}}||{\mathcal{A}}|$).
@@
We are aware that $h\geq\lambda\geq\frac{h^{2}}{2}$, and by polynomially bounding $h$, we can ensure that $\lambda$ is also bounded. In a related work (Liu & Brunskill, 2018), an alternative variation of the Cheeger constant is utilized, which is based on the flow $F$ induced by the stationary distribution $\Psi$ of a random walk on the graph. Suppose for nodes $n_{1},n_{2}$ and subset of nodes $N_{1}$ in the graph, we have:
@@
$F(\sigma N_{1})=\sum_{n_{1}\in N_{1},n_{2}\notin N_{1}}F(n_{1},n_{2}),$ (22)
@@
[…]
5:    $s_{\min}$leftarrow root
@@
11:             Extract option $o$ by actions $root$ to $s_{x^{\prime}}$
@@
14:         if $n(\phi$($s_{x^{\prime}})$) $<$ $n(\phi$($s_{\min})$then
15:             $s_{\min}$=$s_{x^{\prime}}$
@@
19:    Extract option $o$ by actions $root$ to $s_{\min}$
@@
24:    a $\sim$ UniformRandom($\mathcal{A}(s_{x})$)
25:    $s_{x^{\prime}}$leftarrow $\mathcal{T}$($s_{x}$, a)
@@
[…]
@@
Initialize replay buffers $D_{\beta}$, $D_{e}$, hash function $\phi$, exploration budget $N$
@@
     while t $\leq$ T and not(successdo
         if $l$==0 then
              if random()$<\epsilon$ then
                  Exploratory option $w\leftarrow$ generate_option($s$, $\phi$, $N$)
@@
function update($success$)
@@ A.5 Implementation Details and Experimental Hyperparameters
Wall-maze $\mathbb{R}^{2}$ $\mathbb{R}^{2}$ $[-0.95,0.95]^{2}$ 100
U-maze $\mathbb{R}^{6}$ $\mathbb{R}^{2}$ $[-1,1]*[-0.25,0.25]$ 500
Point-push $\mathbb{R}^{11}$ $\mathbb{R}^{2}$ $[-1,1]*[-0.25,0.25]$ 500
window-open $\mathbb{R}^{39}$ $\mathbb{R}^{3}$ $[-1,1]^{4}$ 500
soccer $\mathbb{R}^{39}$ $\mathbb{R}^{3}$ $[-1,1]^{4}$ 500
button-press $\mathbb{R}^{39}$ $\mathbb{R}^{3}$ $[-1,1]^{4}$ 500
@@ A.8 Terminal States Distribution
We analyze the order in which the agent visits different parts of the environment by examining the distribution of the last states in the episodes. To make it more visually appealing and easy to interpret, we only sample some of the episodes. The results for Wall-maze, U-maze, and Point-push can be found in Figures 11, 12, and 13, respectively. In Wall-maze, only $\epsilon t$-greedy and DOIE can effectively navigate to different regions of the environment and ultimately reach the goal area. Other methods often get trapped in one of the local optima and are unable to reach the goal. The reason some methods, such as TD3, have fewer points is that the agent spends a lot of time revisiting congested areas instead of exploring new ones. In U-maze, most methods can explore the majority of the environment. However, during the final stages of training, methods such as DDPG, SAC, and DDPG + intrinsic motivation have lower success rates and may end up in locations other than the goal areas. In Point-push, $\epsilon t$-greedy, $\epsilon z$-greedy, and DOIE first visit the lower section of the environment in the early stages. After that, they push aside the movable box and proceed to the upper section to visit the goal area. For the other methods, the pattern is almost the same, with occasional visits to the lower section.

### v3

@@ A.1 $\epsilon t$-greedy Sample Efficiency : Proofs
We aim to show that the $\epsilon t$-greedy algorithm falls into the PAC-MDP category. According to Liu and Brunskill (2018), an algorithm $\mathcal{A}$ is PAC-MDP if the covering time induced by $\mathcal{A}$ is polynomially bounded. In Liu and Brunskill (2018), the authors further demonstrate that bounding the covering time requires bounding both the Laplacian eigenvalues and the stationary distribution over the states induced by the random walk policy. This is presented as Proposition A.1. According to Theorem 2, two conditions are satisfied: $N\leq\Theta(|\mathcal{S}||\mathcal{A}|)$ and a lower bound on the probability of the sampled option, $\mathcal{P}_{\mathcal{X}}\geq\frac{1}{\Theta(|\mathcal{S}||\mathcal{A}|)}$. These two conditions are necessary and are met by our problem setting and the exploration algorithm (Algorithm 1). To prove that $\mathcal{P}_{\mathcal{X}}\geq\frac{1}{\Theta(|\mathcal{S}||\mathcal{A}|)}$, we construct a worst-case tree structure $\mathcal{X}$, where we aim to identify the option induced by the tree $\mathcal{X}$ with the lowest probability, referred to informally as the “hardest option". We then show that this lower bound satisfies the condition specified in Theorem 1.
@@
[…]
@@
$\Rightarrow\log(N!)+(N-1)c_{0}\leq\log(|{\mathcal{S}}||{\mathcal{A}}|)\quad\quad\quad\text{(Based on the Moivre–Stirling approximation)}$ (12)
@@
First note that if $\mathcal{P}_{\mathcal{X}}[\omega]\geq\frac{1}{\Theta(|\mathcal{S}||\mathcal{A}|)}$ then based on Theorem 1 we need to have $N\leq\frac{\log(|\mathcal{S}||{\mathcal{A}}|)}{\log\log(|\mathcal{S}||{\mathcal{A}}|)}$, and this implies that $N\leq\Theta(|{\mathcal{S}}||{\mathcal{A}}|)$. Based on the paper by (Liu and Brunskill, 2018), and the analysis of the covering length when following a random policy, we have the following preposition:
Preposition A.1 (Liu and Brunskill (2018)).
: For any irreducable MDP M, we define $P_{\pi_{RW}}$ as a transition matrix induced by random walk policy $\pi_{RW}$ over $M$ and $L(P_{\pi_{RW}})$ is denoted as the Laplacian of this transition matrix. Suppose $\lambda$ is the smallest non-zero eigenvalue of $L$ and $\Psi(s)$ is the stationary distribution over states which is induced by random walk policy, then Q-learning with random walk exploration is a PAC RL algorithm if: $\frac{1}{\lambda}$$\frac{1}{\min_{s}\Psi(s)}$ are Poly($|{\mathcal{S}}||{\mathcal{A}}|$).
@@
We are aware that $h\geq\lambda\geq\frac{h^{2}}{2}$, and by polynomially bounding $h$, we can ensure that $\lambda$ is also bounded. In a related work (Liu and Brunskill, 2018), an alternative variation of the Cheeger constant is utilized, which is based on the flow $F$ induced by the stationary distribution $\Psi$ of a random walk on the graph. Suppose for nodes $n_{1},n_{2}$ and subset of nodes $N_{1}$ in the graph, we have:
@@
$F(\sigma N_{1})=\hskip-15.0pt\sum_{n_{1}\in N_{1},n_{2}\notin N_{1}}\hskip-15.0ptF(n_{1},n_{2}),$ (22)
@@
2:  frontier_nodes $\leftarrow\{\}$
3:  Initialize root using $s$: $\mathit{root}\leftarrow\mathit{TreeNode}(s)$
4:  frontier_nodes $\leftarrow$ frontier_nodes $\cup$ {root};
5:  $s_{\min}$ $\leftarrow$ root
@@
11:     Extract option $o$ by actions rootroot to $s_{x^{\prime}}$
@@
14:    if $n(\phi$($s_{x^{\prime}})$) << $n(\phi$($s_{\min})$) then
15:     $s_{\min}$$s_{x^{\prime}}$
@@
19:  Extract option $o$ by actions rootroot to $s_{\min}$
@@
[…]
25:  $s_{x^{\prime}}$ $\leftarrow$ $\mathcal{T}$($s_{x}$, a)
@@
27:  frontier_nodes $\leftarrow$ frontier_nodes $\cup$ $s_{x^{\prime}}$}
@@
Initialize replay buffers $D_{\beta}$, $D_{e}$, hash function $\phi$, exploration budget NN
@@
  while t $\leq$ T and not(success) do
   if ll==0 then
[…]
      Exploratory option $w\leftarrow$ generate_option($s$, $\phi$, NN)
@@
function update(successsuccess)
@@ A.5 Implementation Details and Experimental Hyperparameters
Wall-maze $\mathbb{R}^{2}$ $\mathbb{R}^{2}$ $0.95,0.95]^{2}$ 100
U-maze $\mathbb{R}^{6}$ $\mathbb{R}^{2}$ [1,1][0.25,0.25][-1,1]*[-0.25,0.25] 500
Point-push $\mathbb{R}^{11}$ $\mathbb{R}^{2}$ [1,1][0.25,0.25][-1,1]*[-0.25,0.25] 500
window-open $\mathbb{R}^{39}$ $\mathbb{R}^{3}$ $1,1]^{4}$ 500
soccer $\mathbb{R}^{39}$ $\mathbb{R}^{3}$ $1,1]^{4}$ 500
button-press $\mathbb{R}^{39}$ $\mathbb{R}^{3}$ $1,1]^{4}$ 500
@@ A.8 Terminal States Distribution
We analyze the order in which the agent visits different parts of the environment by examining the distribution of the last states in the episodes. To make it more visually appealing and easy to interpret, we only sample some of the episodes. The results for Wall-maze, U-maze, and Point-push can be found in Figures 11, 12, and 13, respectively. In Wall-maze, only $\epsilon t$-greedy and DOIE can effectively navigate to different regions of the environment and ultimately reach the goal area. Other methods often get trapped in one of the local optima and are unable to reach the goal. The reason some methods, such as TD3, have fewer points is that the agent spends a lot of time revisiting congested areas instead of exploring new ones. In U-maze, most methods can explore the majority of the environment. However, during the final stages of training, methods such as DDPG, SAC, and DDPG + intrinsic motivation have lower success rates and may end up in locations other than the goal areas. In Point-push, $\epsilon t$-greedy, $\epsilon z$-greedy, and DOIE first visit the lower section of the environment in the early stages. After that, they push aside the movable box and proceed to the upper section to visit the goal area. For the other methods, the pattern is almost the same, with occasional visits to the lower section. To further illustrate how the policy evolves during training and how closely it approximates the optimal policy, we present the agent’s trajectories across episodes. For clarity, we sample a few representative episodes from different stages of training. Figure 14 presents the results for U-maze and Point-push. In both environments, the agent initially becomes trapped in suboptimal regions. As training progresses, the policy gradually improves, and by the end, the agent consistently follows a trajectory that closely approximates the shortest path to the goal.
@@
(a) U-maze
(b) Point-push
Figure 14: The agent’s policy during training, illustrated by the paths taken in each episode. Episodes are sampled from different stages of training to show how the policy evolves over time.